I'm having problems on divisibilty on my homework. The exact directions are: Test each number for divisibilty by 2,3,5,9 or 10. Here is ONE problem from my homework. 1) 126 This is the number you have to test. Thanks
Found 2 solutions by ewatrrr, CharlesG2:
Answer by ewatrrr(24785)


Hi,
126
Test for divisibilty by 2,3,5,9 or 10
126/2 = 63
126/3 = 42
126/5 = 25.2
126/9 = 14
126/10 = 12.6
126 is not evenly divisible by 5 or 10

Answer by CharlesG2(834)

I'm having problems on divisibilty on my homework. The exact directions are: Test each number for divisibilty by 2,3,5,9 or 10. Here is ONE problem from my homework. 1) 126 This is the number you have to test. Thanks

126:
by 2, yes since last digit is even
by 3, 1 + 2 + 6 = 9 and 9 is divisible by 3 so yes
by 5, 6 is not divisible by 5 so no
by 9, 126/3 = 42, 4 + 2 = 6 and 6 is divisible by 3 so yes
by 10, the last digit is not 0 so no